What is the wavelength of spectral line resulted from the electron transition from n=3

to n=1 in a hydrogen atom?


1
Expert's answer
2021-12-20T14:23:01-0500

We apply the Rydberg equation;

"\dfrac{1}{\lambda}=Rx(\dfrac{1}{n^2final}-\dfrac{1}{n^2initial})"


"\dfrac{1}{\lambda}=1.0974x10^7m^-1(\dfrac{1}{1^2}-\dfrac{1}{3^2})"


"\dfrac{1}{\lambda}=9754666667"


"{\lambda}=\dfrac{1}{9754666667 }"

"{\lambda} ="1.025 x 10-10 m